OpenStudy (anonymous):

Divide: x^2-9 ---------- x+8 ----------- x^2-6x+9 ----------- x^2+5x-24

OpenStudy (anonymous):

its dividing rational expressions.. i think

OpenStudy (anonymous):

Is the same as =(x^2-9)/((x+8) * (x^2+5x-24)/(x^2-6x+9) =((x-3)(x+3)(x+8)(x-3))/((x+8)(x-3)(x-3)) =x+3

OpenStudy (amistre64):

its like dividing fractions; its easier to just flip the bottom over and multiply

OpenStudy (amistre64):

but you wanna factor it all out after you flip it to see what can cancel

OpenStudy (anonymous):

Eliza's answer n working is correct

OpenStudy (anonymous):

her answer was x+3 correct?

OpenStudy (anonymous):

Yep thanks

OpenStudy (anonymous):

thank you Eliza (: